OpenStudy (anonymous):

In what direction and by how many units is the graph of f(x) = 6 sin(2x + π) − 5 vertically and horizontally shifted? What is the Period?

OpenStudy (dan815):

you want to put it in this form first f(x)= a sin(b(x-c))-d a tells you the amplitude or knows as vertical compression/expansion, b tells you the horizontal compression/expansion, c tells you the horizontal shift, and d tells you the vertical shift

OpenStudy (dan815):

I will show you how to do that for your question, you can try it yourself for some other ones

OpenStudy (dan815):

f(x) = 6 sin(2x + π) − 5 =6 * sin(2*(x-π/2))-5 f(x)= a sin(b(x-c))+d so a=6,b=2,c=π/2,d=-5

OpenStudy (dan815):

a typo here that should be + not - pi/2 so c=-pi/2

OpenStudy (anonymous):

so how do you know how its shifted and what the period is?

OpenStudy (anonymous):

@Rabakku @aaronq @misssunshinexxoxo @ganeshie8 @Hero @jagr2713

OpenStudy (welshfella):

f(x)= a sin(b(x-c))-d f(x) = 6 sin(2x + π) − 5 by comparing the 2 functions you can see that the vertical shift is d = ?